OpenStudy (anonymous):

Solve the equation. Be sure to show your work. All final answers must be in simplest form. f-1/5=1

OpenStudy (anonymous):

F= 6/5

OpenStudy (anonymous):

how did you get f=6/5

OpenStudy (anonymous):

solve for f -> get rid of (-1/5) = ...

OpenStudy (anonymous):

don't get it